# The Hidden Attention of Mamba Models

### v1

@@ 3.3 Application to Attention-based Attribution
As our class-specific explainability technique for Mamba models, we have tailored the Transformer-Attribution [13] explainability method, which is specifically designed for transformers, to suit Mamba models. This method relies on a combination of LRP scores and attention gradients to generate the relevance scores. Since each Mamba block includes several peripheral layers that are not included in transformers, such as Conv1D, additional gating mechanisms, and multiple linear projection layers, a robust mechanism must be carefully designed. For simplicity, we focus on vision Mamba, with a grid of $\sqrt{L}$ patches in each row and column, as in Sec. 3.2.
The Transformer-Attribution method encompasses two stages: (i) generating a relevance map for each attention layer, followed by (ii) the aggregation of these relevance maps across all layers, using the aggregation rule specified in 23, to produce the final map $\rho$.
@@
$\beta^{\lambda}=\mathbb{I}_{L}+\mathop{\mathbb{E}}_{\hat{h}\in[\hat{H}]}(\nabla{\alpha}^{\lambda,\hat{h}}\odot R^{\lambda,\hat{h}})^{+}$ (24)
where $\mathbb{I}_{L+1}\in\mathbb{R}^{(L+1)\times(L+1)}$ is an identity matrix.
Figure 2: Comperative Visualization of Transformer-Attribution (Left), and our Mamba-Attribution (Right).
Our method, Mamba-Attribution, depicted in Fig. 2 (right), deviates from this method by modifying Eq. 24 in the following aspects: (i) Instead of computing the gradients on the per-head attention matrices $\nabla{\alpha}^{\lambda,\hat{h}}$, we compute the gradients of $\nabla\hat{y}^{\lambda,d}$. The motivation for these modifications is to exploit the gradients of both the S6 mixer and the gating mechanism in Eq. 8 (left), to obtain strong class-specific maps. (ii) We simply replace $R^{\lambda,\hat{h}}$ with the attention matrices $\tilde{\alpha}^{\lambda,d}$ at layer $\lambda$ and channel $d$, since we empirically observe that those attention matrices produce better relevance maps. Both of these modifications are manifested by the following form, which defines our method:
$\tilde{\beta}^{\lambda}=\mathbb{I}_{L}+\Big{(}\mathop{\mathbb{E}}_{d\in D}(\nabla\hat{y}^{\lambda,d})\odot\mathop{\mathbb{E}}_{d\in D}(\tilde{\alpha}^{\lambda,d})\Big{)}{}^{+}$ (25)
@@ 4 Experiments
In this section, we present an in-depth analysis of the hidden attention mechanism embedded within Mamba models, focusing on its semantic diversity and applicability in explainable AI frameworks. We start by visualize the hidden attention matrices for both NLP and vision models in Sec. 4.1, followed by empirically assessing our xplainable AI techniques via perturbation and segmentation tests in section 4.2.1.
@@ 4.1 Visualization of Attention Matrices
To better understand the hidden attention mechanism employed in Mamba, as described in Sec. 3.1 and manifested in Eq. 18 and 19, Fig. 3 and 4 containes a comparative visualization of attention matrices in Mamba and Transformer on both vision and NLP tasks. For clearer visualization, we apply the Softmax function to each row of the attention matrices and limit our focus to the initial 64 tokens.
NLP In fig. 3 we compare attention matrices extracted from both Mamba (130m) and Transformer (Pythia-160m [10]) language models, trained on the Pile [20] dataset for next token prediction. The attention maps are extracted using examples from the Lambada dataset (preprocessed by OpenAI).
Figure 3: Hidden Attention Matrices (NLP): The first four columns correspond to the Mamba matrices, while the last column is dedicated to the Transformer. Each row represents a different layer within the models, showcasing the evolution of the attention matrices at 0% (top), 25%, 50%, 75%, and 100% (bottom) of the layer depth.
As can be seen, the hidden attention matrices of Mamba (the first 4 columns) appear similar to the attention matrices extracted from transformers (the last column). In both Mamba and transformers, the dependencies between far-away tokens are captured in the deeper layers of the model, as depicted in the lower rows.
@@
Vision Fig. 4 contains a comparative visualization of attention matrices in Vision-Mamba and ViT (DeiT), for models of tiny size, trained on ImageNet-1K. The attention maps are extracted using examples from the test set. Each Mamba attention matrix is obtained by combining the two maps of the bidirectional channel.
Figure 4: Hidden Attention Matrices (Vision): The first four columns correspond to the Mamba matrices, while the last column is dedicated to the Transformer. Each row represents a different layer within the models, showcasing the evolution of the attention matrices at 0% (top), 25%, 50%, 75%, and 100% (bottom) of the layer depth.
@@ 4.2 Explainability Metrics
The explainable AI experiments include three types of explainability methods: (1) Raw-Attention, which employs raw attention scores as relevancies. Our findings indicate that averaging the attention maps across layers yields optimal results. (2) Attn-Rollout [1] for Transformers, and its Mamba version, as depicted in section 3.2. Finally, (3) The Transformer Attribution of Chefer et al. [12] and its Mamba Attribution counterpart, detailed in Section 3.3.
Fig. 5 depicts the results of the six attribution methods on typical samples from the ImageNet test set. As can be seen, the Mamba-based heatmaps are often more complete than the transformer-based counterparts. The raw attention of Mamba stands out of the other five heatmaps since it is depicts activity across the entire image. However, the relevant object is highlighted.
4.2.1 Pertubation Tests
In this evaluation framework, we employ an input perturbation scheme to assess the efficacy of various explanation methods, following the approach outlined by  [13, 12].
These experiments are conducted under two distinct settings. In the positive perturbation scenario, a quality explanation involves an ordered list of pixels, sorted from the most- to the least-relevant. Consequently, when gradually masking out the pixels of the input image, starting from the highest relevance to the lowest, and measuring the mean top-1 accuracy of the network, one anticipates a notable decrease in performance.
@@ 4.2.1 Pertubation Tests
The results of the perturbations are presented in Table. 1, depicting the performance of different explanation methods under both positive and negative perturbation scenarios across the two models. In the positive perturbation scenario, where lower AUC values are indicative of better performance, we notice that Raw-Attention, Mamba shows a better AUC (17.268 vs. 20.687) compared to the Vision Transformer. For the Attn-Rollout method, Mamba outperforms the Vision Transformer with an AUC of 18.806 vs. 20.594, while the latter shows a better AUC of 15.351 vs. 16.619 under the Attribution method. In the negative perturbation scenario where higher AUC values are better, the Transformer-based methods consistently outperform Mammba across all three methods: Raw-Attention (40.766 vs. 34.025), Attn-Rollout (43.525 vs. 41.864), and Attribution (48.089 vs. 39.632). The tendency for lower AUC in both positive (where it is desirable) and negative perturbation (where it is undesirable) may indicate that the Mamba model is more sensitive to blacking out of patches, and it would be interesting to add experiments in which the patches are blurred instead, following [18].
4.2.2 Segmentation Tests
It is expected that an effective explainability method would produce reasonable foreground segmentation maps. This is assessed for ImageNet classifiers by comparing the obtained heatmap against the ground truth segmentation maps available in the ImageNet-Segmentation dataset [26].
@@ 4.2.2 Segmentation Tests
The results are outlined in Table 2. Raw-Attention, Mamba demonstrates significantly higher pixel accuracy (67.64% vs. 59.69%) and mean Intersection over Union (45.09% vs. 36.94%) compared to Vision Transformer, while the latter performs better in mean Average Precision (77.25% vs. 74.88%). Under the Attn-Rollout method, Mamba outperforms Vision Transformer in mean Average Precision (80.78% vs. 80.34%), pixel accuracy (71.01% vs. 66.84%) and mean Intersection over Union (51.51% vs. 47.85). Finally, Transformer-Attribution consistently surpasses Mamba-Attribution, achieving the highest scores in pixel accuracy (79.26% vs. 74.72%), mean Average Precision (84.85% vs. 81.70%), and mean Intersection over Union (60.63% vs. 54.24%), respectively.
These results underscore the potential of Mamba’s attention mechanism as approaching and sometimes passing the interoperability level of Transformer models, especially when the attention maps are taken as is. It also highlights the applicability of Mamba models for downstream tasks such as weakly supervised segmentation. It seems, however, that the Mamba-based attribution model, which is modeled closely after the transformer method of Chefer et al. [13] may benefit from further adjustments.
@@
Table 1: Positive and Negative perturbation AUC results (percents) for the predicted class on the ImageNet validation set. For positive perturbation lower is better, and for negative perturbation higher is better.
@@ 5 Conclusions
In this work, we have established a significant link between Mamba and self-attention layers, illustrating that the Mamba layer can be reformated as an implicit form of causal self-attention mechanism. This directly links the highly effective Mamba layers with the transformer layers.

### v2

@@ 3.3 Application to Attention-based Attribution
As our class-specific explainability technique for Mamba models, we have tailored the Transformer-Attribution [15] explainability method, which is specifically designed for transformers, to suit Mamba models. This method relies on a combination of LRP scores and attention gradients to generate the relevance scores. Since each Mamba block includes several peripheral layers that are not included in transformers, such as Conv1D, additional gating mechanisms, and multiple linear projection layers, a robust mechanism must be designed carefully. For simplicity, we focus on vision Mamba, with a grid of $\sqrt{L}$ patches in each row and column, as in Sec. 3.2.
The Transformer-Attribution method encompasses two stages: (i) generating a relevance map for each attention layer, followed by (ii) the aggregation of these relevance maps across all layers, using the aggregation rule specified in 21, to produce the final map $\rho$.
@@
$\beta^{\lambda}=\mathbb{I}_{L}+\mathop{\mathbb{E}}_{\hat{h}\in[\hat{H}]}(\nabla{\alpha}^{\lambda,\hat{h}}\odot R^{\lambda,\hat{h}})^{+},\quad\mathbb{I}_{L+1}\in\mathbb{R}^{(L+1)\times(L+1)}$ (22)
Figure 2: Comperative Visualization of Transformer-Attribution and our Mamba-Attribution, both class specific methods.
Figure 3: Average attention maps for CLS token in the middle (a,b,c) and as the first (d,e,f).
Our method, Mamba-Attribution, depicted in Fig. 3, deviates from this method by modifying Eq. 22 in the following aspects: (i) Instead of computing the gradients on the per-head attention matrices $\nabla{\alpha}^{\lambda,\hat{h}}$, we compute the gradients of $\nabla\hat{y}^{\prime\lambda,d}$. The motivation for these modifications is to exploit the gradients of both the S6 mixer and the gating mechanism in Eq. 8 (left), to obtain strong class-specific maps. (ii) We simply replace $R^{\lambda,\hat{h}}$ with the attention matrices $\tilde{\alpha}^{\lambda,d}$ at layer $\lambda$ and channel $d$, since we empirically observe that those attention matrices produce better relevance maps. Both of these modifications are manifested by the following form, which defines our method:
$\tilde{\beta}^{\lambda}=\mathbb{I}_{L}+\Big{(}\mathop{\mathbb{E}}_{d\in D}(\nabla\hat{y}^{\prime\lambda,d})\odot\mathop{\mathbb{E}}_{d\in D}(\tilde{\alpha}^{\lambda,d})\Big{)}{}^{+}$ (23)
@@ 4 Experiments
In this section, we present an in-depth analysis of the hidden attention mechanism embedded within Mamba models, focusing on its semantic diversity and applicability in explainable AI frameworks. We start by visualizing the hidden attention matrices for both NLP and vision models in Sec. 4.1, followed by assessing our explainable AI techniques empirically, via perturbation and segmentation tests in Sec. 4.2.
@@ 4.1 Visualization of Attention Matrices
The Visual Mamba (ViM) comes in two versions: in one, the CLS token is last and in the other, the CLS token is placed in the middle. Fig. 3 shows how this positioning influences the impact of the patches on the CLS, by averaging over the entire test set. Evidently, the patches near the CLS token are more influential. This phenomenon may suggest that a better strategy is to have a non-spatial/global CLS token [22, 37].
Fig. 4 compares the attention matrices in Mamba and Transformer on both vision and NLP tasks. For clearer visualization, we apply the Softmax function to each row of the attention matrices obtained from transformers and perform min-max normalization on the absolute values of the Mamba matrices. In all cases, we limit our focus to the first 64 tokens. In vision, we compare Vision-Mamba (ViM) and ViT (DeiT), for models of a tiny size, trained on ImageNet-1K. The attention maps are extracted using examples from the test set. Each Mamba attention matrix is obtained by combining the two maps of the bidirectional channel. In NLP, we compare attention matrices extracted from Mamba (130m) and Transformer (Pythia-160m [11]) language models, trained on the Pile [25] dataset for next token prediction. The attention maps are extracted using examples from the Lambada dataset (preprocessed by OpenAI).
Vision NLP
Transformer Mamba Transformer Mamba
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Hidden Attention Matrices: Attention matrices in vision and NLP Models.Each row represents a different layer within the models, showcasing the evolution of the attention matrices at 25% (top), 50%, and 75% (bottom) of the layer depth.
As can be seen, the hidden attention matrices of Mamba appear to be similar to the attention matrices extracted from transformers In both models, the dependencies between distant tokens are captured in the deeper layers of the model, as depicted in the lower rows.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d) (e) (f) (g)
@@ 4.2 Explainability Metrics
The explainable AI experiments include three types of explainability methods: (1) Raw-Attention, which employs raw attention scores as relevancies. Our findings indicate that averaging the attention maps across layers yields optimal results. (2) Attn-Rollout [1] for Transformers, and its Mamba version, as depicted in Sec. 3.2. Finally, (3) The Transformer Attribution of Chefer et al. [14] and its Mamba Attribution counterpart, detailed in Sec. 3.3.
Fig. 5 depicts the results of the six attribution methods on typical samples from the ImageNet test set. As can be seen, the Mamba-based heatmaps are often more complete than their transformer-based counterparts. The raw attention of Mamba stands out from the other five heatmaps, since it depicts activity across the entire image. However, the relevant object is highlighted.
Next, we apply explainability evaluation metrics. These metrics allow one to compare different explainability methods that are applied to the same model. Applying them to compare different models is not meant to say that model X is more explainable than model Y. The main purpose is to show that the attention maps of Mamba we introduce are as useful as the attention maps of Transformers in terms of providing explainability. A secondary aim is to validate the feasibility of potential use for weakly supervised downstream tasks that require spatial location. Perturbation
Perturbation Tests  In this evaluation framework, we employ an input perturbation scheme to assess the efficacy of various explanation methods, following the approach outlined by  [15, 14].
These experiments are conducted under two distinct settings. In the positive perturbation scenario, a quality explanation involves an ordered list of pixels, arranged most-to-least relevant. Consequently, when gradually masking out the pixels of the input image, starting from the highest relevance to the lowest, and measuring the mean top-1 accuracy of the network, one anticipates a notable decrease in performance.
@@
The results of the perturbations are presented in Tab. 2, depicting the performance of different explanation methods under both positive and negative perturbation scenarios across the two models. In the positive perturbation scenario, where lower AUC values are indicative of better performance, we notice that for Raw-Attention, Mamba shows a better AUC compared to the Vision Transformer (ViT). For the Attn-Rollout method, Mamba outperforms the ViT, while the latter shows a better AUC under the Attribution method. In the negative perturbation scenario, where higher AUC values are better, the Transformer-based methods consistently outperform Mammba across all three methods. The tendency for lower AUC in both positive (where it is desirable) and negative perturbation (where it is undesirable) may indicate that the Mamba model is more sensitive to blacking out patches, and it would be interesting to add experiments in which the patches are blurred instead [23]. For perturbation tests in the NLP domain, please refer to the Appendices 0.B and 0.C.
Segmentation Tests   It is expected that an effective explainability method would produce reasonable foreground segmentation maps. This is assessed for ImageNet classifiers by comparing the obtained heatmap against the ground truth segmentation maps available in the ImageNet-Segmentation dataset [33].
@@
The results are outlined in Tab. 2. For Raw-Attention, Mamba demonstrates significantly higher pixel accuracy and mean Intersection over Union compared to Vision Transformer, while the latter performs better in mean Average Precision. Under the Attn-Rollout method, Mamba outperforms Vision Transformer in mean Average Precision, pixel accuracy and mean Intersection over Union. Finally, Transformer-Attribution consistently surpasses Mamba-Attribution, achieving the highest scores in pixel accuracy, mean Average Precision, and mean Intersection over Union, respectively.
These results underscore the potential of Mamba’s attention mechanism as approaching and sometimes surpassing the interoperability level of Transformer models, especially when the attention maps are taken as is. It also highlights the applicability of Mamba models for downstream tasks such as weakly supervised segmentation. It seems, however, that the Mamba-based attribution model, which is modeled closely after the transformer method of Chefer et al. [15] may benefit from further adjustments.
Table 1: Positive and Negative perturbation AUC results (percentages) for the predicted class on the ImageNet validation set. For positive perturbation lower is better, and for negative perturbation higher is better.
@@
5 Discussion: The Evolution of Attention in SSMs
A natural question to ask is whether the attention perspective we exposed is unique to Selective SSM (the building block of Mamba), separating it from other SSMs. The answer is that Selective SSM, similar to transformers, contains a type of layer we call data-dependent non-diagonal mixer, which previous layers do not.
In their seminal work, Poli et al. [59] claim that a crucial aspect of transformers is the existence of an expressive, data-controlled linear operator. Here, we focus on a more specific component, which is an expressive data-controlled linear non-diagonal mixer operator. This distinguishes between elementwise operators that act on the data associated with specific tokens (including MLP, gating mechanisms, and GLU activations [65]) and mixer operations that pool information from multiple tokens.
The mixer components can further be divided into fixed, e.g., using pooling operators with fixed structure and coefficients, or data-dependent, in which the interactions between tokens are controlled by their input-dependent representations, e.g., self-attention. In appendix 0.E we prove the following result, which sheds light on the gradual evolution of attention in SSM models.
Theorem 5.1
(i) S4 [31], DSS [34], S5 [66] have fixed mixing elements. (ii) GSS [52],and Hyena [59] have fixed mixing elements with diagonal data-control mechanism. (iii) Selective SSM have data-controlled non-diagonal mixers.
Transformers are recognized for their superior in-context learning (ICL) capabilities, where the model adapts its function according to the input provided [13]. Empirical evidence has demonstrated that Mamba models are the first SSMs to exhibit ICL capabilities on par with those of transformers [29, 56]. Based on the intuition that the capacity to focus on specific inputs is required for ICL, we hypothesize that the presence of data-controlled non-diagonal mixers in both transformers and Mamba models is crucial for achieving a high level of ICL.
A question then arises: which model is more expressive, transformers or selective SSM? While previous work has shown that Transformers are more expressive than traditional state-space layers [85], we show in Appendix 0.D that the situation is reversed for selective SSMs, as follows:
Theorem 5.2
One channel of the selective state-space layer can express all functions that a single transformer head can express. Conversely, a single Transformer layer cannot express all functions that a single selective SSM layer can.
@@ 6 Conclusions
In this work, we have established a significant link between Mamba and self-attention layers, illustrating that the Mamba layer can be reformulated as an implicit form of causal self-attention mechanism. This links the highly effective Mamba layers directly with the transformer layers.